OpenStudy (anonymous):

how did they get this answer: integrate 1/(x^2 + 2x +5)....

OpenStudy (anonymous):

yeh what are you typing?

OpenStudy (anonymous):

\[= 1/((x+1)^2 + 4\] u= x+1 du = 1 dx \[\int\limits_{?}^{?}1/u^2 +4 = 1/2 \tan^{-1} (u/2) +c \] what? how integrate 1/(x^2 +1) but where is the 1/2 coming from?

OpenStudy (anonymous):

it is coming from the 4

OpenStudy (anonymous):

the rest of the problem and that bottom integrate is for arctan \[\int\limits_{?}^{?}1/x^2 +1 = \tan^{-1} x\]

OpenStudy (anonymous):

because 4 = 2^2

OpenStudy (anonymous):

look at this formula sheet http://www.ecalc.com/math-help/worksheet/calculus-integrals

OpenStudy (anonymous):

http://www.ecalc.com/pics/math/calculus_integral.gif look where it says common integrals, and look at the integral 1/ (u^2 + a^2)

OpenStudy (anonymous):

ah im back-- openstudy bonked for a moment --- that equation is a little different than the one in my book...gonna be hard to relearn. And the u's being interchanged ugh. That had to be on purpose.

OpenStudy (anonymous):

how is the one in ur book different?

OpenStudy (anonymous):

It just has: \[\int\limits_{?}^{?}1/x^2 + 1 = \tan^{-1} x + c\]

OpenStudy (anonymous):

yes my bok says integral of du/(u^2 + a^2)=1/a tan^-1 u/a + c

OpenStudy (anonymous):

yes, so the answer makes sense knowing that a= 2 in this case

OpenStudy (anonymous):

4=a^2 therefore a=2

OpenStudy (anonymous):

with the new formula it makes sense...with the old original not so much

OpenStudy (anonymous):

lol ok great!